OpenStudy (anonymous):

what is the standard equation of a sphere centered at origin with point (8,-1,6) on the sphere?

OpenStudy (anonymous):

Find the distance of (8,-1,6) to the origin and that is your radius.

OpenStudy (anonymous):

\[ r = ||(8,-1,6)|| = \sqrt{8^2+1^2+6^2} = \sqrt{101} \]

OpenStudy (anonymous):

So your equation is \[ x^2+y^2+z^2 = 101 \]

OpenStudy (anonymous):

yea, i found it. i was confused because i was thinking of the equation as if it were centered somewhere else at the same time, so i was confused. but now i know. thx! (: